# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v1

@@ 4.1 Statistical Threshold
To this end, we will apply the approach of [PW24] for bipartite graphs, which was inspired by the proof of [Fri90] for ordinary graphs. For each $i\in[r]$, denote the vertices of $V_{i}$ as $\{1_{i},\ldots,n_{i}\}$. Let $m=\left(\dfrac{d}{(\log d)^{r}}\right)^{\frac{1}{r-1}}$, $n^{\prime}=n/m$, and let $P_{i,j}=\{((j-1)m+1)_{i},\ldots,(jm)_{i}\}$ for each $i\in[r]$ and $j\in[n^{\prime}]$. We say $I$ is a $\bm{\gamma}$-balanced $P$-independent set if $I$ is $\bm{\gamma}$-balanced and $|X\cap P_{i,j}|\leq 1$ for each $i\in[r]$ and $j\in[n^{\prime}]$. Let $\beta(H)$ denote the size of the largest $\bm{\gamma}$-balanced $P$-independent set in a hypergraph $H\sim\mathcal{H}(r,n,p)$ and let $Z_{k}$ denote the number of such independent sets of size $k$.
@@ 4.2 Proof of the Achievability Result
In this section, we will construct a degree-$1$ polynomial to prove the first part of Theorem 1.9. In particular, we will construct a polynomial $f\,:\,\{0,1\}^{n^{r}}\to\mathbb{R}^{rn}$ that $(k_{1},\ldots,k_{r},\delta,\xi,0)$-optimizes the maximum $\bm{\gamma}$-balanced independent set problem in $H$ for some $\xi\coloneqq\xi(d,r,\bm{\gamma})\geq 1$ and $\delta=o_{n}(1)$. As mentioned in §1.4, the algorithm we describe is adapted from that of [Dha23] for deterministic hypergraphs.
@@
$k_{i}=\gamma_{i}\,r\,n\,(1-\varepsilon)\left(\frac{\log d}{d\,r^{r}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r.$
If some $k_{i}$ is smaller, we can zero out some coordinates to achieve this. Furthermore, assume $\varepsilon\leq 1/(100r)$ and $\gamma_{r}=\max_{i\in[r]}\gamma_{i}$. In particular, $\gamma_{r}\geq 1/r$.
@@
$\geq\exp\left(-\frac{(1-\varepsilon)^{r-1}}{(1-r\varepsilon/2)}\cdot\frac{1}{r(r-1)\gamma_{r}}\cdot\log d\right)$
@@
where we use the fact that $\gamma_{r}r\geq 1$ and let $\tilde{\varepsilon}=r\varepsilon/4$. Let $\mu\coloneqq\mathbb{E}\left[\sum_{v\in V_{r}}f_{v}(A)\right]$. In particular, we may conclude that
$\frac{k_{r}}{\mu}\leq r\,\gamma_{r}\,\left(\frac{\log d}{d^{\tilde{\varepsilon}}\,r^{r}\,(r-1)\,\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\leq\frac{1}{2},$
@@ 4.3 Forbidden Structures
Fix $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Let $K\in\mathbb{N}$ be a sufficiently large constant in terms of $r$, $\varepsilon$, and $\bm{\gamma}$. Consider an interpolation path $A^{(0)},\ldots,A^{(T)}$ of length $T=n^{O(r)}$. If $d\coloneqq d(\varepsilon,r,\bm{\gamma},K)$ is sufficiently large, then with probability $1-\exp\left(-\Omega(n)\right)$ there does not exist a sequence of subsets $S_{1},\ldots,S_{K}$ of $V$ satisfying the following:
@@
    for each $k\in[K]$ and $i\in[r]$, we have $|S_{k}\cap V_{i}|\geq(1+\varepsilon)\,\gamma_{i}\,r\,n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}$.
@@
    $|S_{k}\setminus\bigcup_{j<k}S_{j}|\in\left[\frac{\tilde{\varepsilon}}{4}\,n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\,\frac{\tilde{\varepsilon}}{2}\,n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\right].$
@@
Let $N$ denote the number of tuples $(S_{1},\ldots,S_{K})$ satisfying conditions (B1)(B3). We will show that $\mathbb{E}[N]$ is exponentially small (the result then follows by Markov’s inequality). To this end, let
$\Phi\coloneqq n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},$
@@
Conditions (B2) and (B3) imply that $a_{i,k}\geq(1+\varepsilon)\gamma_{i}r$ and $b_{k}\in[\tilde{\varepsilon}/4,\,\tilde{\varepsilon}/2]$. Then for $c=|\cup_{k}S_{k}|/\Phi$, we have $c\leq\sum_{i}a_{i,k}+(K-1)\tilde{\varepsilon}/2$. As a result of the upper bound in Theorem 1.8, we may assume that $\sum_{i}a_{i,k}\leq(1+\varepsilon)r^{r/(r-1)}$. In particular, $c=C(r,\varepsilon,K)$ is a bounded constant independent of $d$. There are at most $n^{2rK}$ choices for $\{a_{i,k}\}$ and $\{b_{i,k}\}$ ($\{b_{k}\}$ can be determined from $\{b_{i,k}\}$). Once these are fixed, the number of tuples $(S_{1},\ldots,S_{K})$ is at most
@@
$=\frac{\Phi\,n^{r-1}\log d}{d(r-1)}\left(\prod_{i=1}^{r}\frac{a_{i,1}}{r\gamma_{i}}+\sum_{k=2}^{K}\left(\prod_{i=1}^{r}\frac{a_{i,k}}{r\gamma_{i}}-\prod_{i=1}^{r}\frac{(a_{i,k}-b_{i,k})}{r\gamma_{i}}\right)\right).$
@@
$|E|\geq\frac{\Phi\,n^{r-1}\log d}{d(r-1)}\left(\sum_{i=1}^{r}\tilde{a}_{i,k}-r+1+\sum_{k=2}^{K}(1+\varepsilon/2)^{r-1}\sum_{i=1}^{r}\tilde{b}_{i,k}\right).$
@@
$\qquad\exp\left(-\frac{\Phi\log d}{r-1}\left(\sum_{i=1}^{r}\tilde{a}_{i,k}-r+1+\sum_{k=2}^{K}(1+\varepsilon/2)^{r-1}\sum_{i=1}^{r}\tilde{b}_{i,k}\right)\right)$
$\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\,\log d}{r-1}\left(r-1+\sum_{i=1}^{r}(a_{i,1}-\tilde{a}_{i,1})\right)\right)$
$\qquad\exp\left(\frac{\Phi\log d}{r-1}\left(\sum_{k=2}^{K}\left(b_{k}-(1+\varepsilon/2)^{r-1}\sum_{i=1}^{r}\tilde{b}_{i,k}\right)+o_{n,d}(1)\right)\right).$
Note the following:
$\tilde{b}_{i,k}=\frac{b_{i,k}}{r\gamma_{i}}\geq\frac{b_{i,k}}{r\max_{j}\gamma_{j}}.$
In particular, for $\varepsilon$ small enough, we have the following for some constant $\alpha\coloneqq\alpha(\varepsilon,\bm{\gamma})>0$:
$\mathbb{E}[N]$ $\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\,\log d}{r-1}\left(r-1+\sum_{i=1}^{r}a_{i,1}-\alpha\sum_{k=2}^{K}b_{k}+o_{n,d}(1)\right)\right)$
Once again, as a result of Theorem 1.8, we may assume that $\sum_{i=1}^{r}a_{i,1}\leq(1+\varepsilon)r^{r/(r-1)}$. As $b_{k}\geq\tilde{\varepsilon}/4$, for $K$ sufficiently large in terms of $r$, $\varepsilon$, and $\bm{\gamma}$, we have
@@
Fix $\varepsilon,s>0$ and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Let $d=d(\varepsilon,s,\bm{\gamma})>0$ be sufficiently large, and for each $i\in[r]$ fix any $S_{i}\subseteq V_{i}$ such that
$|S_{i}|\,\leq\,s\,\gamma_{i}\,r\,n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}.$
@@
$|S^{\prime}\cap S_{i}|\,\geq\,\varepsilon\,\gamma_{i}\,r\,n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
Let $\Phi$ be as defined in the proof of Proposition 4.4. We will compute the expected number of such sets $S^{\prime}$ and show it is exponentially small (the proof then follows by Markov’s inequality). Let $N$ denote the number of possible $S^{\prime}$. We have
@@
$\leq\left(\prod_{i=1}^{r}\left(\frac{es}{\varepsilon}\right)^{\varepsilon\gamma_{i}r\Phi}\right)\exp\left(-\frac{\varepsilon^{r}\,\Phi\log d}{r-1}\right)$
$=\exp\left(-\Phi\left(\frac{\varepsilon^{r}\,\log d}{r-1}-\varepsilon\,r\,\log\left(\frac{es}{\varepsilon}\right)\right)\right)$
@@ 4.4 Proof of the Intractability Result
For any $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$, there exist $K,\,d_{0}>0$ such that for any $d\geq d_{0}$ there exists $n_{0},\eta,C_{1},C_{2}>0$ such that for any $n\geq n_{0}$, $\xi\geq 1$, $1\leq D\leq\dfrac{C_{1}n}{\xi\log n}$, and $\delta\leq\exp\left(-C_{2}\xi\,D\log n\right)$, if
$k_{i}\,\geq\,(1+\varepsilon)\,\gamma_{i}\,r\,n\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
Fix $r$, $\varepsilon$, and $\bm{\gamma}$ as described in the statement of the proposition. We define the following parameters:
$\Phi=\left(\frac{\log d}{d\,r^{r}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\qquad\tilde{\varepsilon}=\varepsilon\min_{i\in[r]}\gamma_{i}.$
@@
$(1+\varepsilon)\gamma_{i}\,r\,n\,\Phi\,\leq\,|I_{t}\cap V_{i}|\,\leq\,(1+\varepsilon)\,r^{r/(r-1)}\,n\,\Phi,\qquad\text{for all }0\leq t\leq T,\,\,1\leq i\leq r,$
@@
$S_{i}=S\cap V_{i},\quad\text{and}\quad s=\frac{(1+\varepsilon)\,r^{r/(r-1)}K^{\prime}}{\min_{i}\gamma_{i}},$

### v2

@@ 4.1 Statistical Threshold
To this end, we will apply the approach of [perkins2024hardness] for bipartite graphs, which was inspired by the proof of [frieze1990independence] for ordinary graphs. For each $i\in[r]$, denote the vertices of $V_{i}$ as $\{1_{i},\ldots,n_{i}\}$. Let $m=\left(\dfrac{d}{(\log d)^{r}}\right)^{\frac{1}{r-1}}$, $n^{\prime}=n/m$, and let $P_{i,j}=\{((j-1)m+1)_{i},\ldots,(jm)_{i}\}$ for each $i\in[r]$ and $j\in[n^{\prime}]$. We say $I$ is a $\bm{\gamma}$-balanced $P$-independent set if $I$ is $\bm{\gamma}$-balanced and $|X\cap P_{i,j}|\leq 1$ for each $i\in[r]$ and $j\in[n^{\prime}]$. Let $\beta(H)$ denote the size of the largest $\bm{\gamma}$-balanced $P$-independent set in a hypergraph $H\sim\mathcal{H}(r,n,p)$ and let $Z_{k}$ denote the number of such independent sets of size $k$.
@@ 4.2 Proof of the Achievability Result
In this section, we will construct a degree-$1$ polynomial to prove the first part of Theorem 1.9. In particular, we will construct a polynomial $f\,:\,\{0,1\}^{n^{r}}\to\mathbb{R}^{rn}$ that $(k_{1},\ldots,k_{r},\delta,\xi,0)$-optimizes the maximum $\bm{\gamma}$-balanced independent set problem in $H$ for some $\xi\coloneqq\xi(d,r,\bm{\gamma})\geq 1$ and $\delta=o_{n}(1)$. As mentioned in §1.4, the algorithm we describe is adapted from that of [dhawan2023balanced] for deterministic hypergraphs.
@@
$k_{i}=\gamma_{i}\,r\,n\,(1-\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r.$
If some $k_{i}$ is smaller, we can zero out some coordinates to achieve this. Furthermore, assume $\varepsilon\leq 1/(100r)$ and $i^{\star}=r$.
@@
$\geq\exp\left(-\frac{(1-\varepsilon)^{r-1}}{(1-r\varepsilon/2)}\cdot\frac{1}{r-1}\cdot\log d\right)$
@@
where we let $\tilde{\varepsilon}=r\varepsilon/4$. Let $\mu\coloneqq\mathbb{E}\left[\sum_{v\in V_{r}}f_{v}(A)\right]$. In particular, we may conclude that
$\frac{k_{r}}{\mu}\leq r\,\gamma_{r}\,\left(\frac{\gamma_{i^{\star}}\log d}{d^{\tilde{\varepsilon}}\,r^{r-1}\,(r-1)\,\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\leq\frac{1}{2},$
@@ 4.3 Forbidden Structures
Fix $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. Let $K\in\mathbb{N}$ be a sufficiently large constant in terms of $r$, $\varepsilon$, and $\bm{\gamma}$. Consider an interpolation path $A^{(0)},\ldots,A^{(T)}$ of length $T=n^{O(r)}$. If $d\coloneqq d(\varepsilon,r,\bm{\gamma},K)$ is sufficiently large, then with probability $1-\exp\left(-\Omega(n)\right)$ there does not exist a sequence of subsets $S_{1},\ldots,S_{K}$ of $V$ satisfying the following:
@@
    for each $k\in[K]$ and $i\in[r]$, we have $|S_{k}\cap V_{i}|\geq(1+\varepsilon)\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}$.
@@
    $|S_{k}\setminus\bigcup_{j<k}S_{j}|\in\left[\frac{\tilde{\varepsilon}}{4}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\,\frac{\tilde{\varepsilon}}{2}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\right].$
@@
Without loss of generality, let $i^{\star}=r$. Let $N$ denote the number of tuples $(S_{1},\ldots,S_{K})$ satisfying conditions (B1)(B3). We will show that $\mathbb{E}[N]$ is exponentially small (the result then follows by Markov’s inequality). To this end, let
$\Phi\coloneqq n\left(\frac{\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r-1}\gamma_{i}}\right)^{1/(r-1)},$
@@
Conditions (B2) and (B3) imply that $a_{i,k}\geq(1+\varepsilon)\gamma_{i}r$ and $b_{k}\in[\tilde{\varepsilon}/4,\,\tilde{\varepsilon}/2]$. Then for $c=|\cup_{k}S_{k}|/\Phi$, we have $c\leq\sum_{i}a_{i,k}+(K-1)\tilde{\varepsilon}/2$. As a result of the upper bound in Theorem 1.8, we may assume that $\sum_{i}a_{i,k}\leq(1+\varepsilon)r\,\gamma_{r}^{-1/(r-1)}$. In particular, $c=C(r,\varepsilon,K)$ is a bounded constant independent of $d$. There are at most $n^{2rK}$ choices for $\{a_{i,k}\}$ and $\{b_{i,k}\}$ ($\{b_{k}\}$ can be determined from $\{b_{i,k}\}$). Once these are fixed, the number of tuples $(S_{1},\ldots,S_{K})$ is at most
@@
$=\frac{\Phi\,r\gamma_{r}\,n^{r-1}\log d}{d(r-1)}\left(\prod_{i=1}^{r}\frac{a_{i,1}}{r\gamma_{i}}+\sum_{k=2}^{K}\left(\prod_{i=1}^{r}\frac{a_{i,k}}{r\gamma_{i}}-\prod_{i=1}^{r}\frac{(a_{i,k}-b_{i,k})}{r\gamma_{i}}\right)\right).$
@@
$|E|$ $\geq\frac{\Phi\,r\gamma_{r}\,n^{r-1}\log d}{d(r-1)}\left(\sum_{i=1}^{r}\tilde{a}_{i,k}-r+1+\sum_{k=2}^{K}(1+\varepsilon/2)^{r-1}\sum_{i=1}^{r}\tilde{b}_{i,k}\right)$
$\geq\frac{\Phi\,n^{r-1}\log d}{d(r-1)}\left(\sum_{i=1}^{r}a_{i,k}-r\gamma_{r}(r-1)+\sum_{k=2}^{K}(1+\varepsilon/2)^{r-1}\sum_{i=1}^{r}b_{i,k}\right)$
where we use the fact that $\gamma_{r}\geq\gamma_{i}$ for each $i$ by definition of $i^{\star}$.
@@
$\qquad\exp\left(-\frac{\Phi\log d}{r-1}\left(\sum_{i=1}^{r}a_{i,k}-r\gamma_{r}(r-1)+\sum_{k=2}^{K}(1+\varepsilon/2)^{r-1}b_{k}\right)\right)$
$\leq n^{2rK}(T+1)^{k}\sup_{\{a_{i,k}\},\{b_{i,k}\}}\exp\left(\frac{\Phi\log d}{r-1}\left(r\gamma_{r}(r-1)-\frac{r\varepsilon}{4}\sum_{k=2}^{K}b_{k}+o_{n,d}(1)\right)\right).$
As $b_{k}\geq\tilde{\varepsilon}/4$, for $K$ sufficiently large in terms of $r$ and $\varepsilon$, we have
@@
Fix $\varepsilon,s>0$ and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. Let $d=d(\varepsilon,s,\bm{\gamma})>0$ be sufficiently large, and for each $i\in[r]$ fix any $S_{i}\subseteq V_{i}$ such that
$|S_{i}|\,\leq\,s\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}.$
@@
$|S^{\prime}\cap S_{i}|\,\geq\,\varepsilon\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
Without loss of generality, assume $i^{\star}=r$. Let $\Phi$ be as defined in the proof of Proposition 4.4. We will compute the expected number of such sets $S^{\prime}$ and show it is exponentially small (the proof then follows by Markov’s inequality). Let $N$ denote the number of possible $S^{\prime}$. We have
@@
$\leq\left(\prod_{i=1}^{r}\left(\frac{es}{\varepsilon}\right)^{\varepsilon\gamma_{i}r\Phi}\right)\exp\left(-\frac{\varepsilon^{r}\,\gamma_{r}r\,\Phi\log d}{r-1}\right)$
$=\exp\left(-\Phi\left(\frac{\varepsilon^{r}\,\gamma_{r}r\,\log d}{r-1}-\varepsilon\,r\,\log\left(\frac{es}{\varepsilon}\right)\right)\right)$
@@ 4.4 Proof of the Intractability Result
For any $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$, $\sum_{i=1}^{r}\gamma_{i}=1$, and $i^{\star}=\arg\max_{i}\gamma_{i}$, there exist $K,\,d_{0}>0$ such that for any $d\geq d_{0}$ there exists $n_{0},\eta,C_{1},C_{2}>0$ such that for any $n\geq n_{0}$, $\xi\geq 1$, $1\leq D\leq\dfrac{C_{1}n}{\xi\log n}$, and $\delta\leq\exp\left(-C_{2}\xi\,D\log n\right)$, if
$k_{i}\,\geq\,(1+\varepsilon)\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
Fix $r$, $\varepsilon$, and $\bm{\gamma}$ as described in the statement of the proposition. Additionally, assume with out loss of generality that $i^{\star}=r$. We define the following parameters:
$\Phi=\left(\frac{\gamma_{r}\,\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\qquad\tilde{\varepsilon}=\varepsilon\min_{i\in[r]}\gamma_{i}.$
@@
$(1+\varepsilon)\gamma_{i}\,r\,n\,\Phi\,\leq\,|I_{t}\cap V_{i}|\,\leq\,(1+\varepsilon)\,r\gamma_{r}^{-1/(r-1)}\,n\,\Phi,\qquad\text{for all }0\leq t\leq T,\,\,1\leq i\leq r,$
@@
$S_{i}=S\cap V_{i},\quad\text{and}\quad s=\frac{(1+\varepsilon)\,r\gamma_{r}^{-1/(r-1)}K^{\prime}}{\min_{i}\gamma_{i}},$